1. the grams of water that must be added to 4.00 g KI in the preparation of 2.00 % KI solution.


1
Expert's answer
2021-10-27T06:46:23-0400

Let the mass of solvent be x.

"Percentage\space concentration=2=\frac{2}{x+2}×100" "2x+4=200\space \space \space x=\frac{200-4}{2}=98grams"